# Forecasting Events in Soccer Matches Through Language

### v1

@@ Abstract.
This paper introduces an approach to predicting the next event in a soccer match, a challenge bearing remarkable similarities to the problem faced by Large Language Models (LLMs). Unlike other methods that severely limit event dynamics in soccer, often abstracting from many variables or relying on a mix of sequential models, our research proposes a novel technique inspired by the methodologies used in LLMs. These models predict a complete chain of variables that compose an event, significantly simplifying the construction of Large Event Models (LEMs) for soccer. Utilizing deep learning on the publicly available WyScout dataset, the proposed approach notably surpasses the performance of previous LEM proposals in critical areas, such as the prediction accuracy of the next event type. This paper highlights the utility of LEMs in various applications, including betting and match analytics. Moreover, we show that LEMs provide a simulation backbone on which many analytics pipelines can be built, an approach opposite to the current specialized single-purpose models. LEMs represent a pivotal advancement in soccer analytics, establishing a foundational framework for multifaceted analytics pipelines through a singular machine-learning model.
Soccer Event Prediction, Large Event Models, Deep Learning, Simulation, Sports Analytics
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Information systems Data miningccs: Computing methodologies Discrete-event simulation
Figure 1. Types of events occurring during a professional soccer game. The figure depicts a soccer field with various symbols representing events during a professional soccer game (Real Madrid - Barcelona, December 23, 2017). The types of events include passes, duels, shots, and fouls. Other event types not included in the figure are goalkeeper saves, cards, free kicks, among others.
1. Introduction
@@
Soccer, despite its global popularity and economic impact, has lagged behind other sports in leveraging data analytics for insights. This gap stems from the inherent dynamism of the game, exceeding other team sports due to the larger number of players and constant interactions. As a result, soccer presents a vast, fertile ground for advancements in sports analytics, particularly through leveraging the power of artificial intelligence.
Recent developments in AI, specifically the emergence of foundational models capable of learning the full spectrum of a game and facilitating insightful interpretation, offer exciting possibilities for revolutionizing soccer analytics. Large Event Models (LEMs) (Mendes-Neves et al., 2024) exemplify this paradigm by presenting a novel approach to generating soccer events, serving as a valuable tool for developing generative models in this domain. Still, LEMs have several limitations.
@@
We present a novel approach to building LEMs that addresses limitations of existing LEMs. Instead of relying on sequential inferences from multiple models, we draw inspiration from the core methodologies of language models. By tokenizing soccer event data, we enable a single model to effectively learn the ”language” of soccer events. This single model performs sequential inferences, with each inference corresponding to a token representing a part of an event. Our model captures the most relevant aspects of soccer events with seven inferences, achieving superior performance in forecasting key variables compared to conventional LEMs.
@@
  • Betting: LEMs can predict the likelihood of future events, empowering users to make informed betting decisions based on a comprehensive understanding of game dynamics.
@@
2. Related Work
Soccer has witnessed a growing trend in data-driven analysis, aiming to uncover patterns and predict future outcomes. From injury prediction (Kakavas et al., 2020) to transfer values (McHale and Holmes, 2023), there are several applications where machine learning can help soccer teams and players improve their performance.
However, these applications are driven through specialized models. For each new application, a new model needs to be built from scratch, which is time costly and prone to errors. To solve this, simulation environments (Kurach et al., 2020) and approaches (Mendes-Neves et al., 2021) have been proposed. However, they are still limited in scope.
LEMs show promise in regards to progress towards a general approach (Mendes-Neves et al., 2024). It shows significant progresses over previous approaches (Simpson et al., 2022) which were limited in terms of scope, mostly due to only using a small portion of offensive events. By (1) extending the covered events and (2) significantly accelerating the inference process, the original LEM proposal enabled large-scale simulations of soccer matches. These simulations can be used for many use cases, including those stated before injury prediction, valuing player contribution, forecasting probabilities, among others.
@@ 2. Related Work
  1. (1)
    Complex Architecture: LEMs use three different deep learning models to predict different aspects of a soccer event.
  2. (2)
@@
LLMs can predict complete sentences: they do not need a model to forecast nouns and another to forecast adjectives. The ability of LLMs to learn general knowledge and generate coherent data inspired the exploration of approaches similar to LLMs with the aim of solving the problem of forecasting the next event in a soccer match.
The application of techniques used in LLMs to other areas is not something novel. LLMs have transcended text generation, demonstrating success in diverse domains like music composition (van den Oord et al., 2016) and physical simulation (Jain et al., 2022).
Generative artificial intelligence is still in the early stages for soccer. Our proposal provides a ground foundation to advance LEMs toward real applications in soccer by providing a significant improvement in terms of model accuracy and ease of implementation.
3. A Language-based Approach to LEMs
3.1. Data
We use the Public Wyscout dataset, which contains data from Wyscout V2 API (Pappalardo et al., 2019). The dataset contains data for the 5 most valuable leagues in European football: England, Spain, Germany, Italy, and France. This dataset does not represent the most up-to-date data standards available in the industry. However, publicly available datasets with the most up-to-date data. For each event, we extract the key features highlighted in Figure 2.
We use the France, Germany, and Italy leagues on our training set, leaving the complete seasons of England and Spain for testing. Although not including data from the testing league in the training set leads to worse performance, we opted to leave these leagues completely out of the dataset so they could be used for season-long application development using our models, without the risk of overfitting.
@@ 3.1. Data
  • Event Type: Categorical variable indicating of the event, e.g, pass, shot, take on, tackle.
@@
  • isHomeTeam: Binary variable indicating if the action was performed by the home team.
@@
3.2. Leveraging Language Models for Event Prediction
Soccer can be abstracted to Markov chains (Rudd, 2011; Decroos et al., 2019). A soccer match is a sequence of events performed by the players on the field. This sequential nature of soccer makes it a fertile ground for Markov-based approaches to forecast many aspects of the game, e.g., next event (Simpson et al., 2022; Mendes-Neves et al., 2024), action value (Decroos et al., 2019). It is easy to draw the parallel between generating the next events and generating the next word. As LLMs use the previous words as context to forecast the next word, LEMs use the previous events as context to forecast the next event.
There are several aspects of LLMs which are relevant for our approach. LLMs provide a basis for sequencial generation, which is crucial - we do not want to forecast the next event, but rather a chain of future events and their outcomes. Sequential generation is what enables LEMs to generate full soccer matches, enabling the extraction of multiple metrics from the data. LLMs are also very good at understanding previous context. Context in soccer is key. The next event should reflect the current state of the game, including several long-term aspects such as current score.
In abstraction, LLMs tokenize and encode the text to facilitate the learning problem. Tokenization is the process of breaking down a large piece of text or data into smaller units called tokens (Otter et al., 2021). For example, the word “learning” can be split in the tokens “learn” and “ing”. This facilitates the algorithm’s learning process in multiple aspects. The “ing” token is repeated across multiple instances in the English language - decoupling the “ing” from “learning” will enable the reutilization of the same token across all gerund words. Then, the tokens need to be encoded into numbers, which can be used by deep learning models.
For our proposal, we do not perform tokenization of event data. The vocabulary in event data is extremely limited, with only 140 different words used for the selected data points. There is no semantic aspect to our data, other than sequenciality. Therefore, in this use case, tokenization does not have a significant impact on the performance of our models.
The same rational can be applied to the encoding. The simplicity of our approach makes advanced approaches, like embeddings (Otter et al., 2021), inefficient. We employed an ordinal encoding approach, where tokens are encoded in the following way:
@@ 3.2. Leveraging Language Models for Event Prediction
  • categorical values regarding action types occupy the next 37 positions of the encoder, ordered from the most to least frequent.
  • ¡PERIOD_OVER¿ and ¡GAME_OVER¿ tokens occupy the next 2 positions of the encoder.
  • the ¡NaN¿ token occupies the last position of the encoder.
@@
Table 1. This table shows a sample of the raw event data extracted from the Wyscout dataset.
@@
Table 2. This table presents the encoded version of the event data from Table 1, using ordinal encoding. This format is suitable for machine learning algorithms.
Our approach fixes several limitations of previous proposals: (1) it simulates all action types, overcoming the limitation of only being able to simulate the attacking portion of the game (Simpson et al., 2022), (2) it generates events from a singular model, severely simplifying the architecture and preserving complete sequentiality of the generation process.
@@
3.3. Formulation
Figure 2. A schematic representation of our proposal. In blue, we have the set of inputs used to build the input vector, which is passed through an LEM to infer the probabilities of each token. To make a prediction, the probabilities go through a samples with restrictions to avoid hallucinations.
@@ 3.3. Formulation
The predicted $\hat{p}$ vector is then used to sample the next token in the event, through a Multinomial sampler. The sampler contains restrictions regarding which part of the event is being predicted. The restrictions act as a diversity control system. If it is predicting the first token within an event, corresponding to the event type, then only the probabilities related to event type tokens are passed to the sampler. The restrictions are required to ensure consistent outputs from the models. As the model will be used for millions of inferences and simulations, any residual probability that is not consistent with the event data format can derail the prediction of future events, making the simulation unusable.
@@
But our approach can only forecast one token at the time. To solve this problem, we need to provide the model with information about the current composition of the event. The model needs to be informed about which position of the event tupple is being predicted. To achieve this, the predicted elements are added to the set $S$. This way, the model is informed about which positions of the event were already predicted and can then better forecast the next position within the event. For modeling purposes, we pad the flattened vector with the ¡NaN¿ token, which the model can use to determine which predictions are valid for the next token.
@@
There is two exceptions in the prediction of the next event which is not the same as the input:
@@
  • Current scores for home and away teams are computed externally, avoiding the burden of extra inferences to calculate a variable which is deterministically computable.
@@
  • The Period, Minute and Second variables are transformed into a TimeElapsed variable. This shortens the number of inferences required by two. All three variables are then deterministically computed from the previous values of the variables, the predicted Event Type and Time Elapsed variables.
3.4. Deep Learning
@@
Recent advances in deep learning architectures, such as RNNs (Hochreiter and Schmidhuber, 1997) and Transformers (Radford et al., 2018), enabled large LLM performance increases. However, these architectures increase the number of learnable parameters in the neural networks. In LLMs, the increase in the number of parameters can be easily compensated with an increase in computational power and larger datasets. GPT-2 models (Radford et al., 2019) use 40gb of data to train a 1.5B parameter model.
@@
Due to the constrains in publicly available soccer event data, we opted to use a multi-layer perceptron architecture for our model. The available event data in our dataset, after pre-processed, is 100mb. Although we see significant gains to be made using more advanced architectures, the size of data available to us is not sufficient to thoroughly explore these approaches.
Our architecture was determined empirically, detailed as follows:
@@
  • Hidden Layers: The model contains 3 layers with 512 neurons each, with the lite version of the model containing 2 layers with 256 neurons each. For the k=1 model, these architectures contain around 600k and 100k parameters respectively. We use ReLU as the activation function.
@@
  • Learning Rate: Initiated at a value of 0.001, the learning rate undergoes dynamic adjustment each epoch, following a cosine annealing schedule (Huang et al., 2017). This approach aids in stabilizing the learning process over epochs.
@@
  • Optimizer: The Adam optimizer (Kingma and Ba, 2017) was employed for its efficiency in handling sparse gradients and adaptive learning rate capabilities.
One of the LEMs’ main purposes is to enable the large-scale simulation of soccer games. For such, the user only needs to iteratively query the model, while arranging the current state of the game to be updated by the new predictions.
Moreover, the LEM is implemented in such way that the only limiting factor for the number of parallel simulations is the amount of available GPU memory. This architecture allows the parallel simulation of millions of soccer matches concurrently, enabling simulation-based insights at large scale.
In some games, the current event has frequent long-term impacts on the outcome of matches. Due to soccer’s low scoring nature, there are few factors which can significantly impact the current game status, completely changing the context of the match. This is a great feature for our models - as soccer contains a lot of “reset points”, such as set pieces, or whenever the ball goes out-of-bound, a new sequence is started. With the start of a new sequence, the cumulative error of our predictions has a soft reset. While some variables, such as the current score, have a significant impact in the following outcome, these resets provide our model with an opportunity to rebalance its predictions in cases where the model starts to halucinate.
3.5. Implementation

### v2

@@ Abstract
This paper introduces an approach to predicting the next event in a soccer match, a challenge bearing remarkable similarities to the problem faced by Large Language Models (LLMs). Unlike other methods that severely limit event dynamics in soccer, often abstracting from many variables or relying on a mix of sequential models, our research proposes a novel technique inspired by the methodologies used in LLMs. These models predict a complete chain of variables that compose an event, significantly simplifying the construction of Large Event Models (LEMs) for soccer. Utilizing deep learning on the publicly available WyScout dataset, the proposed approach notably surpasses the performance of previous LEM proposals in critical areas, such as the prediction accuracy of the next event type. This paper highlights the utility of LEMs in various applications, including match prediction and analytics. Moreover, we show that LEMs provide a simulation backbone for users to build many analytics pipelines, an approach opposite to the current specialized single-purpose models. LEMs represent a pivotal advancement in soccer analytics, establishing a foundational framework for multifaceted analytics pipelines through a singular machine-learning model.
[…]
2331
1 Introduction
@@
Despite its global popularity and economic impact, soccer has lagged behind other sports in leveraging data analytics for insights. This gap stems from the inherent dynamism of the game, exceeding other team sports due to the larger number of players and constant interactions. As a result, soccer presents a vast, fertile ground for advancements in sports analytics, mainly through leveraging the power of artificial intelligence.
Recent developments in AI, specifically the emergence of foundational models capable of learning the full spectrum of a game and facilitating insightful interpretation, offer exciting possibilities for revolutionizing soccer analytics. Large Events Models (LEMs) [11] exemplify this paradigm. LEMs represent a novel approach in the field of sports analytics, particularly tailored for understanding and predicting sequences of events in soccer. Unlike traditional models that rely on isolated statistical metrics, LEMs draw inspiration from the architecture of Large Language Models (LLMs), which have demonstrated remarkable success in natural language processing. By conceptualizing a soccer match as a series of events — each akin to a "word" in a sentence — LEMs utilize deep learning techniques to predict the next event in the sequence based on the context provided by previous events. This methodology allows for a dynamic and contextually rich analysis of sports games, capturing events’ fluid and unpredictable nature as they unfold during a match.
The core innovation of LEMs lies in their ability to predict discrete outcomes and generate probabilistic forecasts for every aspect of an event, enabling the generation of entire sequences of game events and offering a comprehensive tool for game simulation and analysis. This ability is achieved through sequential deep learning models, each responsible for forecasting a specific aspect of the event. By feeding data into a series of predictive models, LEMs can simulate various possible future scenarios of a match from any given moment, thereby providing insights into likely game developments. Still, LEMs have several limitations.
@@
We present a novel approach to building LEMs that addresses the limitations of existing LEMs, depicted in Figure 1. Instead of relying on sequential inferences from multiple models, we draw inspiration from the core methodologies of language models. By tokenizing soccer event data, we enable a single model to learn the "language" of soccer events effectively. This single model performs sequential inferences, with each inference corresponding to a token representing a part of an event. Our model captures the most relevant aspects of soccer events with seven inferences, achieving superior performance in forecasting key variables compared to conventional LEMs.
Figure 1: A schematic representation of our proposal. In blue, we have the set of inputs used to build the input vector, passed through the LEM model to infer the probabilities of each token. To make a prediction, the probabilities go through a sampler with restrictions to avoid hallucinations, i.e., predicting unrealistic values.
@@
  • Match Prediction: LEMs can predict the likelihood of future events, empowering users to make informed decisions based on a comprehensive understanding of game dynamics.
@@
2 Related Work
Soccer has witnessed a growing trend in data-driven analysis, aiming to uncover patterns and predict future outcomes. From injury prediction [5] to transfer values [8], there are several applications where machine learning can help soccer teams and players improve their performance.
However, these applications are driven through specialized models. For each new application, a new model needs to be built from scratch, which is time-consuming and prone to errors. To solve this, simulation environments [7] and approaches [10] have been proposed. However, they are still limited in scope.
LEMs [11] draw inspiration from Large Language Models (LLMs), which are trained on vast datasets and can be fine-tuned for specific tasks. By treating sequences of game events as "sentences" of actions, where the model predicts the next "word" describing the event based on the "context" of previous events, LEMs provide a generative framework to make evaluations in soccer. As soccer matches are sequential, this framework allows the simulation of soccer matches from start to end. LEMs employ a sequential forecasting method to align and predict multiple variables more accurately, predicting variables in a staged approach as follows:
  • Model Type: This stage predicts the type of the next event.
  • Model Accuracy: This stage assesses the accuracy of the upcoming event and predicts whether it will result in a goal.
  • Model Data: This final stage predicts the location of the next event, the time until this event occurs, and whether the home team will be involved.
In each stage, the LEM generates a probability distribution for each variable. For example, the output from the Model Type provides probabilities for each potential event type. Similar to the approach used in LLMs, our model samples from these probabilities to determine the next event type. Therefore, each forecast is probabilistic, implying that the output might vary for the same contextual input based on the predicted probabilities.
LEMs show promise regarding progress towards a general approach [11]. It shows significant progress over previous approaches [17], which were limited in scope, mostly due to only using a small portion of offensive events. By (1) extending the covered events and (2) significantly accelerating the inference process, the original LEM proposal enabled large-scale simulations of soccer matches. These simulations can be used for many use cases, including those stated before injury prediction, valuing player contribution, and forecasting probabilities.
@@ 2 Related Work
  1. 1.
    Complex Architecture: LEMs use three different multi-layer perceptron models to predict different aspects of a soccer event. One model predicts the event type, then a different model predicts the accuracy of the event, and only then is the model able to infer spatiotemporal aspects of the event.
  2. 2.
@@
The approach proposed in this paper provides a substantial improvement for all of these drawbacks: (1) We use a single model to forecast the whole event, which simplifies architecture, removes synchronization challenges, and reduces the number of hyperparameters requiring tuning, and (2) each aspect of the event is calculated using the full context of the previous forecasts, increasing the amount of information, and consequently leading to an increase in performance.
LLMs can predict complete sentences: they do not need a model to forecast nouns and another to forecast adjectives. The ability of LLMs to learn general knowledge and generate coherent data inspired the exploration of approaches similar to LLMs to solve the problem of forecasting the next event in a soccer match.
Applying techniques used in LLMs to other areas is not novel. LLMs have transcended text generation, demonstrating success in diverse domains like music composition [18] and physical simulation [4].
Generative artificial intelligence is still in the early stages of soccer. Our proposal provides a ground foundation to advance LEMs toward real applications in soccer by providing a significant improvement in terms of model accuracy and ease of implementation.
3 A Language-based Approach to LEMs
3.1 Data
We use the Public Wyscout dataset, which contains data from Wyscout V2 API [13]. The dataset includes data for the 2017/18 season of the five most valuable leagues in European football: England, Spain, Germany, Italy, and France. This dataset does not represent the most up-to-date data standards available in the industry, but it is one of the most up-to-date publicly available datasets. We considered merging datasets from different sources. However, the standards between providers have substantial differences, which would incur more limitations to our work. We extract the key features highlighted in Figure 1 for each event.
We use the France, Germany, and Italy leagues on our training set, leaving the complete seasons of England and Spain for testing. Although not including data from the testing league in the training set leads to worse performance, we opted to leave these leagues entirely out of the dataset so they could be used for season-long application development using our models without the risk of overfitting. This limitation is exclusive of the publicly available datasets and should
@@ 3.1 Data
  • Event Type: Categorical variable indicating the event, e.g, pass, shot, take on, tackle.
@@
  • isHomeTeam: Binary variable indicating if the home team performed the action.
@@
Some of the features could be merged since they are co-dependent: movement across the x coordinate affects movement across the y coordinate. However, combining these two variables would lead to a large increase in our output space (50+ times larger) and, consequently, a worse computational performance. Therefore, the variables are predicted sequentially, with the x variable being predicted first due to its higher importance relative to the y coordinate.
The model does not include features like team and player identification since they substantially increase the input and output space. There are over 2000 players in our dataset, increasing the model size. Including specific team or player information data on LEMs remains an area for future research.
@@
3.2 Leveraging Language Models for Event Prediction
Soccer can be abstracted to Markov chains [16, 1]. A soccer match is a sequence of events performed by the players on the field. This sequential nature of soccer makes it a fertile ground for Markov-based approaches to forecasting many aspects of the game, e.g., next event [17, 11], action value [1]. It is easy to draw the parallel between generating the next events and generating the next word. As LLMs use the previous words as context to forecast the next word, LEMs use the previous events as context to forecast the next event.
Several aspects of LLMs are relevant to our approach. LLMs provide a basis for sequential generation, which is crucial - we do not want to forecast the next event, but rather a chain of future events and their outcomes. Sequential generation enables LEMs to generate full soccer matches, enabling the extraction of multiple metrics from the data. LLMs are also very good at understanding previous contexts. Context in soccer is key. The next event should reflect the game’s current state, including several long-term aspects such as the current score.
In abstraction, LLMs tokenize and encode the text to facilitate the learning problem. Tokenization is breaking down a large text or data into smaller units called tokens [12]. For example, the word "learning" can be split into the tokens "learn" and "ing". This step facilitates the algorithm’s learning process in multiple aspects. The "ing" token is repeated across multiple instances in the English language - decoupling the "ing" from "learning" will enable the reutilization of the same token across all gerund words. Then, the tokens need to be encoded into numbers, which deep learning models can use.
For our proposal, we do not perform tokenization of event data. The vocabulary in event data is minimal, with only 140 words used for the selected data points. There is no semantic aspect to our data other than sequentially. Therefore, in this use case, tokenization does not significantly impact our models’ performance.
The same rationale can be applied to the encoding. The simplicity of our approach makes advanced approaches, like embeddings [12], inefficient. Employing an ordinal encoding approach is more practical. The tokens are encoded in the following way:
@@ 3.2 Leveraging Language Models for Event Prediction
  • categorical values regarding action types occupy the following 37 encoder positions, from the most to least frequent.
  • <PERIOD_OVER> and <GAME_OVER> tokens occupy the following two encoder positions.
  • the <NaN> token occupies the last position of the encoder.
@@
Table 1: This table shows a sample of the raw event data extracted from the Wyscout dataset.
@@
Table 2: This table presents the encoded version of the event data from Table 1, using ordinal encoding. This format is suitable for machine learning algorithms.
Our approach fixes several limitations of previous proposals: (1) it simulates all action types, overcoming the limitation of only being able to simulate the attacking portion of the game [17], (2) it generates events from a singular model, severely simplifying the architecture and preserving complete sequentiality of the generation process.
@@
3.3 Formulation
@@
The predicted $\hat{p}$ vector is then used to sample the next event token through a Multinomial sampler. The sampler contains restrictions regarding which part of the event is being predicted. The restrictions act as a diversity control system. If it predicts the first token within an event, corresponding to the event type, then only the probabilities related to event type tokens are passed to the sampler. The restrictions are required to ensure consistent outputs from the models. As the model will be used for millions of inferences and simulations, any residual probability inconsistent with the event data format can derail the prediction of future events, making the simulation unusable.
@@
However, our approach can only forecast one token at a time. To solve this problem, we need to provide the model with information about the current composition of the event. The model needs to be informed about which position of the event tuple is being predicted. The predicted elements are added to the set $S$ to achieve this. This way, the model is informed about which event positions were already predicted and can better forecast the next position within the event. For modeling purposes, we pad the flattened vector with the <NaN> token, which the model can use to determine which predictions are valid for the next token.
@@
There are two exceptions in the prediction of the next event, which is not the same as the input:
@@
  • Current scores for home and away teams are computed externally, avoiding the burden of extra inferences to calculate a deterministically computable variable.
@@
  • The Period, Minute, and Second variables are transformed into a TimeElapsed variable. This transformation shortens the number of inferences required by two. All three variables are then deterministically computed from the previous values of the variables, the predicted Event Type, and Time Elapsed variables.
3.4 Deep Learning
@@
Recent advances in deep learning architectures enabled large LLM performance increases, such as RNNs [2] and Transformers [14]. However, these architectures substantially increase the number of learnable parameters in the neural networks. In LLMs, the increase in the number of parameters can be easily compensated with an increase in computational power and larger datasets. GPT-2 models [15] use 40GB of data to train a 1.5B parameter model.
@@
Due to the constraints in publicly available soccer event data, we used a multi-layer perceptron architecture for our model. After pre-processing, the available event data in our dataset is 100 MB. Although we see significant gains to be made using more advanced architectures, the size of the data available to us is insufficient to explore these approaches thoroughly.
Our architecture was determined by trial and error. We increased the size of the model layers in powers of 2 until we found no tangible gain over the previous iteration. We repeated the process for 2 and 3 layers. The final architectures are detailed as follows:
@@
  • Hidden Layers: The model contains three layers with 512 neurons each, with the lite model containing two layers with 256 neurons each. For the k=1 model, these architectures contain around 600k and 100k parameters, respectively. We use ReLU as the activation function. Although the architecture seems overparameterized when compared with the number of features available, it is important to note that we have a much larger output space.
@@
  • Learning Rate: Initiated at a value of 0.001, the learning rate undergoes dynamic adjustment each epoch, following a cosine annealing schedule [3]. This approach aids in stabilizing the learning process over epochs.
@@
  • Optimizer: The Adam optimizer [6] was employed for its efficiency in handling sparse gradients and adaptive learning rate capabilities.
One of the LEMs’ main purposes is to enable the large-scale simulation of soccer games. For such, the user only needs to iteratively query the model, while arranging the game’s current state to be updated by the new predictions.
Moreover, the LEM is implemented so that the only limiting factor for the number of parallel simulations is the amount of available GPU memory. This architecture allows the parallel simulation of millions of soccer matches concurrently, enabling simulation-based insights at a large scale.
In some games, the current event has frequent long-term impacts on the outcome of matches. Due to soccer’s low-scoring nature, a few factors can significantly impact the current game status, completely changing the context of the match. This is a great feature for our models - as soccer contains a lot of "reset points", such as set pieces, or whenever the ball goes out-of-bound, a new sequence is started. With the start of a new sequence, the cumulative error of our predictions has a soft reset. While some variables, such as the current score, have a significant impact on the following outcome, these resets allow our model to rebalance its predictions in cases where the model starts to hallucinate.
3.5 Implementation and Reproducibility